# Phonon-Coupled Hole-Spin Qubits in High-Purity Germanium: Design and Modeling of a Scalable Architecture

### v2

@@ VI Device Design Concept
Figure 11 further illustrates this core qubit structure. A single hole is electrostatically confined near the interface of a strained Ge quantum well and a high-$\kappa$ gate dielectric. The surrounding gate architecture enables lateral and vertical confinement, while a co-integrated Pncc supports coherent spin-phonon interactions for control and coupling.
Figure 11: Schematic illustration of a Ge-based hole-spin qubit integrated with a PnCC. The qubit is electrostatically confined by gate electrodes at the interface between the cap layer and a strained Ge quantum well, grown on a relaxed Si1-xGex buffer. The epitaxial strain enhances spin–orbit coupling and coherence. A lateral PnCC is positioned adjacent to the qubit to enable coupling between the localized hole spin and engineered acoustic phonon modes. The cavity acts as a spectral filter and mediator for coherent spin manipulation and long-range qubit interactions. The phononic mode is spatially aligned with the quantum dot to maximize spin–phonon coupling efficiency.
@@
The entire system operates at millikelvin temperatures in a dilution refrigerator, where spin initialization is achieved via gate pulsing, and control is implemented through EDSR. Readout is performed through spin-to-charge conversion using quantum point contacts (QPCs) or single-electron transistors (SETs) placed adjacent to the quantum dot [49, 8]. This detection method offers high-fidelity spin measurement and fast response suitable for scalable architectures.
@@ VII.1 Simulation Models and Equations
where $g_{0}=2.0$ is the zero-field $g$-factor and $\alpha=0.7\,\mathrm{(MV/m)}^{-1}$ is the tunability coefficient. Simultaneously, the spin-phonon coupling strength $g_{sp}$ increases linearly with the applied field:
@@
where the constant $C$ is chosen such that $T_{1}=1\,\mathrm{ms}$ at $f=6\,\mathrm{GHz}$, yielding $C=36\,\mathrm{ms\cdot GHz^{2}}$.
@@ VII.2 Quantum Dot Behavior and Spin-Orbit Coupling
Compared to other quantum dot platforms, such as GaAs electron spins where the g-factor is fixed near −$-0.44$ [4], or Si spin qubits which exhibit weak spin-orbit coupling [5], Ge hole-spin qubits offer superior electrical tunability. This enables fast EDSR control without the need for additional micromagnets or microwave antennas [7, 8].
@@ VII.3 Phonon Dynamics: Relaxation and Confinement
The magnitude of $M_{\text{sp}}$ depends on several key parameters, including the spatial overlap between the phonon-induced strain and the qubit wavefunction, the degree of heavy-hole/light-hole mixing, the applied electric field (which modulates Rashba spin–orbit interaction), and the orientation of the external magnetic field relative to the crystallographic axes [11, 8]. In our simulation framework, $M_{\text{sp}}$ is numerically computed by integrating the product of the simulated phononic strain profile and the modeled wavefunction envelope over the dot region. This methodology provides realistic, geometry-aware estimates of the transition matrix element that controls both $T_{1}$ relaxation and phonon-mediated qubit dynamics.
@@
where $\omega$ is the mode frequency and $\Delta\omega$ the linewidth obtained from eigenmode simulations.
@@
These results compare favorably with other solid-state systems. For instance, superconducting qubits coupled to piezoelectric resonators exhibit shorter $T_{1}\sim\mu$s due to substrate losses [55]. In contrast, our Ge-based design achieves millisecond-scale lifetimes and high Q-factors within engineered bandgaps, without relying on piezoelectric effects or optical coupling [37, 46].
@@ VII.5 Experimental Validation Strategies
    Spin Readout: Employ QPC or SET-based charge sensors for single-shot readout via spin-to-charge conversion. These sensors also enable detection of phonon-assisted tunneling.
@@ VII.6 Design Benchmark Comparison
To assess the relevance and performance of the proposed Ge-based phonon-coupled qubit architecture, we compare our simulated results to design targets established in prior experimental and theoretical works. These targets span g-factor tunability, spin-phonon coupling strength, relaxation times, and phonon cavity Q-factors—key metrics determining qubit operability and scalability. Tables 3–5 summarize this comparison.
@@
Spin-phonon coupling $>$5 MHz [10, 11] 0.5–6.3 MHz
@@
T1 relaxation time $>$0.5 ms [11, 28] 0.6–5.0 ms
Q-factor $>$104 [37, 46] 10k–18k
@@
[…]
@@ VIII Discussion
To address these fabrication-related issues, it is critical to implement careful optimization of etching parameters, post-etch surface passivation techniques, and controlled thermal annealing processes. Moreover, to effectively mitigate strain-induced disorder, the use of lattice-matched relaxed Si1-xGex buffer layers (with $x\approx 0.95$–$1.0$) or isotopically enriched 74Ge substrates is recommended. Ongoing experimental validation efforts will be essential to quantify the practical impacts of these imperfections on device performance. Insights gained from such experiments will inform subsequent design optimizations, ensuring robust and reliable operation of Ge-based quantum devices under realistic fabrication conditions.

### v3

@@ VI Device Design Concept
Figure 11 further illustrates this core qubit structure. A single hole is electrostatically confined near the interface of a strained Ge quantum well and a high-$\kappa$ gate dielectric. The surrounding gate architecture enables lateral and vertical confinement, while a co-integrated Pncc supports coherent spin-phonon interactions for control and coupling. We consider two integration schemes for the phononic-crystal cavity (PnCC) that preserve the relaxed Si1-xGex buffer and the two-dimensional hole gas (2DHG): (i) a suspended cap-level membrane above the quantum well (QW), and (ii) an under-etched device layer beneath the QW. In both designs, the charge-sensor channel is lithographically isolated from the suspended region, and the confined acoustic mode couples to the dot via evanescent strain, so hole confinement and sensor electrostatics are not degraded. The PnCC membrane is laterally offset from the charge sensor and vertically separated from the 2DHG by a spacer to preserve the local electrostatics; arrows indicate the displacement field of the confined mode.
Figure 11: Schematic illustration of a Ge-based hole-spin qubit integrated with a PnCC. Electrostatic gates confine a hole at the cap/strained-Ge quantum-well interface grown on a relaxed Si1-xGex buffer; the epitaxial strain enhances spin–orbit coupling and coherence. A lateral PnCC adjacent to the dot supports a localized GHz mechanical mode inside an engineered bandgap, enabling controlled coupling between the localized spin and acoustic phonons. The cavity serves as a spectral filter and mediator for coherent spin manipulation and long-range qubit interactions; the quantum dot is positioned at a displacement antinode (yellow in the schematic) to maximize the spin–phonon coupling $g_{\rm sp}$. Inset: A second, identical dot at separation $d$ couples via the confined mode; tuning yields an effective XY interaction $H_{\rm int}=\tfrac{\hbar J}{2}\!\left(\sigma_{x}^{(1)}\sigma_{x}^{(2)}+\sigma_{y}^{(1)}\sigma_{y}^{(2)}\right)$ in the dispersive regime. The rightward arrow indicates energy flow along the localized cavity mode.
@@
The entire system operates at millikelvin temperatures in a dilution refrigerator, where spin initialization is achieved via gate pulsing, and control is implemented through EDSR. Readout is performed via two complementary approaches: (i) dispersive phonon-cavity readout using a $6~\mathrm{GHz}$ mode ($Q=10\mathrm{k}\text{--}18\mathrm{k}$; $\kappa/2\pi\approx 0.33\text{--}0.60~\mathrm{MHz}$), which, together with $g_{\mathrm{sp}}\approx 6~\mathrm{MHz}$ and $\Delta/2\pi=10\text{--}50~\mathrm{MHz}$, yields single-shot $\tau\approx 2\text{--}10~\mu\mathrm{s}$ with $98\%$ fidelity; and (ii) spin-to-charge conversion detected by a proximal QPC/SET positioned outside the PnCC defect and acoustically isolated by shallow trenches/phononic shield. The proximal QPC/SET charge sensor is located outside the PnCC defect and couples to the quantum dot [59, 8] through an electrostatic channel. Acoustic isolation trenches (and optional phononic shields for suspended membranes) inhibit energy leakage from confined mode, preventing $Q$ degradation while allowing standard spin-to-charge single-shot readout. This detection method offers high-fidelity spin measurement and fast response suitable for scalable architectures.
@@ VII.1 Simulation Models and Equations
where $g_{0}$ and $\alpha$ are obtained by linear regression to six-band $\mathbf{k}\!\cdot\!\mathbf{p}$ results with Bir–Pikus strain over $E_{z}\in[0,1.0]~\mathrm{MV/m}$, $g_{0}=2.0$ is the zero-field $g$-factor and $\alpha=0.7\,\mathrm{(MV/m)}^{-1}$ is the tunability coefficient. Simultaneously, the spin-phonon coupling strength $g_{sp}$ increases linearly with the applied field:
@@
(phenomenological fit over 2–6 GHz to capture the observed trend that higher-frequency modes yield shorter $T_{1}$); see Refs. [44, 45, 46] for microscopic derivations and alternative scalings, where the constant $C$ is chosen such that $T_{1}=1\,\mathrm{ms}$ at $f=6\,\mathrm{GHz}$, yielding $C=36\,\mathrm{ms\cdot GHz^{2}}$.
@@ VII.2 Quantum Dot Behavior and Spin-Orbit Coupling
Compared to other quantum dot platforms, such as GaAs electron spins where the g-factor is fixed near 0.44-0.44 [4], or Si spin qubits which exhibit weak spin-orbit coupling [5], Ge hole-spin qubits offer superior electrical tunability. This enables fast EDSR control without the need for additional micromagnets or microwave antennas [7, 8].
@@ VII.3 Phonon Dynamics: Relaxation and Confinement
The magnitude of $M_{\text{sp}}$ depends on several key parameters, including the spatial overlap between the phonon-induced strain and the qubit wavefunction, the degree of heavy-hole/light-hole mixing, the applied electric field (which modulates Rashba spin–orbit interaction), and the orientation of the external magnetic field relative to the crystallographic axes [11, 8]. In our simulation framework, $M_{\rm sp}$ is computed from the Bir–Pikus spin–phonon Hamiltonian $H_{\rm so}[u]$ as a matrix element between qubit states, and $g_{\rm sp}=M_{\rm sp}/\hbar$. In the acoustic regime $H_{\rm so}[u]\propto\varepsilon(\mathbf{r})$, which motivates the overlap formulation. This methodology provides realistic, geometry-aware estimates of the transition matrix element that controls both $T_{1}$ relaxation and phonon-mediated qubit dynamics.
@@
[…]
@@
Given our simulated spin-phonon coupling strengths in the range of $g_{sp}=5\text{–}10~\text{MHz}$ and assuming a qubit-cavity detuning of $\Delta=100~\text{MHz}$, the resulting inter-qubit coupling strength is $g_{qq}\sim 0.25\text{–}1~\text{MHz}$. This value is sufficient to enable two-qubit gate operations on sub-microsecond timescales—well within the $T_{2}^{*}$ coherence times observed in Ge spin qubits (typically $\sim 1\text{–}10~\mu\text{s}$[8, 59], ensuring high-fidelity phonon-mediated entanglement in a realistic experimental setting.
@@ VII.5 Experimental Validation Strategies
  • Spin Readout: Employ QPC or SET-based charge sensors for single-shot readout via spin-to-charge conversion. These sensors also enable detection of phonon-assisted tunneling. Here ‘phonon-assisted tunneling’ denotes inelastic tunneling between adjacent dots or between a dot and a reservoir, accompanied by absorption/emission of a confined phonon; such events modulate the sensor signal in dispersive readout.
@@ VII.6 Design Benchmark Comparison
To assess the relevance and performance of the proposed Ge-based phonon-coupled qubit architecture, we compare our simulated results to design targets established in prior experimental and theoretical works. These targets span g-factor tunability, spin-phonon coupling strength, relaxation times, and phonon cavity Q-factors—key metrics determining qubit operability and scalability. Tables 35 summarize this comparison.
@@
Spin-phonon coupling >>5 MHz [10, 11] 0.5–6.3 MHz
@@
T1 relaxation time >>0.5 ms [11, 30] 0.6–5.0 ms
Q-factor >>104 [39, 56] 10k–18k
@@
The benchmarking results summarized in Tables V–VII demonstrate that the proposed phonon-coupled Ge hole-spin qubit architecture achieves spin–phonon coupling strengths and quality factors that are competitive with or exceed those of current EDSR-based qubit platforms. For comparison, recent experiments in planar Si/Ge heterostructures have reported electrically driven single-hole spin qubits with Rabi frequencies exceeding $100~\text{MHz}$ and coherence times $T_{2}^{*}\sim 1\text{--}5~\mu\text{s}$ [8, 31]. These systems rely on strong Rashba spin–orbit coupling and local gate control for fast manipulation but remain limited in terms of scalable qubit–qubit coupling mechanisms and noise isolation from broadband phonons.
@@ VIII Discussion
To address these fabrication-related issues, it is critical to implement careful optimization of etching parameters, post-etch surface passivation techniques, and controlled thermal annealing processes. Moreover, to effectively mitigate strain-induced disorder, the use of lattice-matched relaxed Si1-xGex buffer layers (with $x\approx 0.95$$1.0$) or isotopically enriched 74Ge substrates is recommended. Ongoing experimental validation efforts will be essential to quantify the practical impacts of these imperfections on device performance. Insights gained from such experiments will inform subsequent design optimizations, ensuring robust and reliable operation of Ge-based quantum devices under realistic fabrication conditions.